OpenStudy (anonymous):

Find the period of the minute hand of a wall clock. Find the angular frequency of the minute hand of a wall clock. Find the frequency of the minute hand of a wall clock.

OpenStudy (anonymous):

i can tell you that the period is 3600s, if you can also explain this i would be greatful

OpenStudy (ganpat):

1. Period = 60 min * 60 = 3600 Sec 2. Angular velocity = (2 * Pi * N ) 60.. = ( 2 * pi * 360) / 60 = 37.699 . 3. Frequency = 1/N = 1/ 360 ... I guess, i am right... @Vaidehi09 : what say ? u good at Physics..

OpenStudy (anonymous):

it says that they are incorrect. thanks for trying.

OpenStudy (ganpat):

lol, m sorry !!!

OpenStudy (anonymous):

Period T is right. but angular velocity w = 2 pi /T = 2pi/3600 = 91/1800 rad/sec and frequency = 1/T = 1/3600

OpenStudy (anonymous):

it says angular velocity is 1.75*10^-3. is your answer in rad/s?

OpenStudy (ganpat):

yes ..

OpenStudy (anonymous):

the frequency was 1/3600 thank you for that.

OpenStudy (anonymous):

can anyone explain angular velocity please? and thanks for the effort ganpat

OpenStudy (anonymous):

hey wait.....there was a typing error....w = pi/1800 rad/s

OpenStudy (anonymous):

aww bummer, too bad i only had one try left on that question. thanks though now that i understand it better

OpenStudy (anonymous):

glad to help :) just remember the formulae and it'll be a piece of cake!